Circular No. 8056

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

XTE J1720–318 C. B. Markwardt, University of Maryland and Goddard Space Flight

Center (GSFC); and J. H. Swank, GSFC, write: “The best x-ray position of the transient XTE J1720–318, derived from RXTE PCA scans on Jan. 15.06 UT, is α = 17^{h}20^{m}00^{s}, δ = −31^{o}44′.8 (equinox 2000.0). The estimated systematic uncertainty is 0′.5, consistent with the radio position (IAUC 8054). The x-ray flux declined 30 percent between Jan. 14 and 19, from a maximum of 1.0 × 10^{−8} erg cm^{−2} s^{−1} (2–10 keV). The soft spectrum and low high-frequency variability are consistent with the high soft states of black-hole candidates.”

D. Kato and T. Nagata, Nagoya University, on behalf of the InfraRed Survey Facility (IRSF)/Simultaneous three-color InfraRed Imager for Unbiased Survey (SIRIUS) team of Nagoya University and the National Astronomical Observatory of Japan; T. Kato and M. Uemura, Kyoto University; and H. Yamaoka, Kyushu University, report on near-infrared (J, H, K_{s}) imaging of the field of the x-ray transient XTE J1720–318 (IAUC 8050) with the IRSF (which consists of the Nagoya-South African Astronomical Observatory 1.4-m telescope + SIRIUS) at Sutherland on Jan. 18.1 and 19.1 UT. A likely infrared counterpart was detected at α = 17^{h}19^{m}59^{s}.00, δ = −31^{o}45′01′′.2 (equinox 2000.0; ∼ 150 2MASS reference stars; rms error 0′′.2), located within 1′′.7 of the proposed radio counterpart (IAUC 8054). The new object is invisible on public 2MASS images; any definite identification would require further follow-up observations, since the brightness of the new object is close to the limiting magnitudes of the 2MASS survey. Magnitudes for the likely infrared counterpart of XTE J1720–318 are: Jan. 18.1, J = 16.73 ± 0.02, H = 15.91 ± 0.03, K_{s} = 15.31 ± 0.03; 19.1, 16.68 ± 0.02, 15.80 ± 0.03, 15.27 ± 0.05.

SUPERNOVA 2002ka W. M. Wood-Vasey, Lawrence Berkeley National Laboratory, reports

that prediscovery NEAT observations of SN 2002ka (cf. IAUC 8038), taken with the NEAT 1.2-m reflector at Haleakala, yield unfiltered magnitudes of 15.7 on 2002 Dec. 13.7 UT and 15.8 on Dec. 19.1.

SUPERNOVA 2002jy IN NGC 477 Unfiltered CCD magnitude estimates by T. Vanmunster, Landen, Bel-

gium: 2002 Dec. 26.817 UT, 15.9; 2003 Jan. 8.733, 16.6.

2003 January 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green